# TeLL-Drive: Enhancing Autonomous Driving with Teacher LLM-Guided Deep Reinforcement Learning

### v1

@@ V-C1 Comparison with Baseline Methods
TABLE I: Comparison of security, efficiency, and real-time test results of different methods in multiple scenarios.
@@
In unsignalized intersection shown in Fig. 5(a), Our model rapidly improves during the initial training phase and converges to the highest final return. A2C exhibits significant fluctuations, implying instability near convergence. While other baselines eventually stabilize, their end-stage returns remain notably below ours, highlighting a substantial performance gap. As shown in Fig. 5(b), our method achieves high returns early on, stabilizing around $5\times 10^{4}$ steps and consistently maintaining near-optimal performance thereafter. In contrast, DQN starts with negative returns and steadily climbs to a suboptimal plateau. A2C fares the worst, likely owing to its sensitivity in time-critical merging tasks. Although RecurrentPPO converges more promptly than A2C, its ultimate reward remains below our model’s, underscoring the challenges of handling highly dynamic traffic with simple recurrent mechanisms. All methods experience rapid early gains, yet differ significantly in final returns and stability in Fig. 5(c). Our model maintains a leading position throughout and converges to a near-maximal reward. RecurrentPPO is intermittently competitive but prone to fluctuations. DQN and A2C both show moderate terminal performance, with A2C stabilizing late but still achieving a lower reward ceiling.
What’s more, Table I provides a numerical summary of success rate, evaluation return, average speed, $\Delta$TTCP, and decision-making time for each approach. In the unsignalized intersection scenario, our method attains the highest success rate (88%) while balancing speed and safety margins. For high-speed ramp merging, it achieves 91% success and outperforms the baselines in average return. Notably, A2C, despite having the highest speed, completely fails (0% success), demonstrating that overly aggressive driving sacrifices safety and thus overall performance. In four-lane adaptive cruise, our method reaches a perfect success rate (100%) alongside near-optimal speed and return. Although Dilu shows better speed and safety margins than traditional DRL methods, its extended reasoning time limits online deployment. Overall, our approach surpasses both conventional DRL algorithms and the LLM-based Dilu, underlining its effectiveness and robustness across diverse scenarios.
V-C2 Ablation Study
V-C3 Teacher-Student Comparison
V-D Case Analysis
@@
To illustrate the model’s decision-making and behavioral patterns in realistic settings, Fig. 8 presents three representative cases. In Fig. 8(a), the agent carefully negotiates an unsignalized intersection, slowing at step 4 to yield and again decelerating at step 12 upon detecting oncoming traffic. Fig. 8(b) shows the agent merging onto a highway by accelerating, then briefly braking to maintain a safe following distance, and finally speeding up once the gap is sufficient. In the denser traffic scenario of Fig. 8(c), the agent continuously adapts its speed to preserve a safe headway despite flow disturbances. Taken together, these examples corroborate our quantitative results, evidencing the method’s efficacy in real-time decision-making, traffic interaction, and safety.
VI Conclusion
@@
Our proposed TeLL-Drive framework integrates teacher-guided learning with attention-based policy optimization, enabling efficient knowledge transfer and robust decision-making. Experimental results demonstrate that TeLL-Drive outperforms conventional DRL methods and existing LLM-based approaches across multiple metrics, including success rate, average return, and real-time feasibility. Additionally, ablation studies highlight the significance of each model component, particularly the synergy between attention mechanisms and LLM teacher guidance. These findings confirm that our approach not only accelerates policy convergence but also enhances safety and adaptability across diverse traffic conditions. In the future, we will explore the application of the TeLL-Drive framework to more dynamic, multi-agent environments and verify its scalability and real-time adaptability through real-vehicle experiments.

### v3

@@ V-C1 Comparison with Baseline Methods
In unsignalized intersection shown in Fig. 5(a), Our model rapidly improves during the initial training phase and converges to the highest final return. A2C exhibits significant fluctuations, implying instability near convergence. While other baselines eventually stabilize, their end-stage returns remain notably below ours, highlighting a substantial performance gap. In high-speed ramp merging shown in Fig. 5(b), our method achieves high returns early on, stabilizing around $5\times 10^{4}$ steps and consistently maintaining near-optimal performance thereafter. In contrast, DQN starts with negative returns and steadily climbs to a suboptimal plateau. A2C fares the worst, likely owing to its sensitivity in time-critical merging tasks. Although RecurrentPPO converges more promptly than A2C, its ultimate reward remains below our model’s, underscoring the challenges of handling highly dynamic traffic with simple recurrent mechanisms. All methods experience rapid early gains, yet differ significantly in final returns and stability in four-lane adaptive cruise shown in Fig. 5(c). Our model maintains a leading position throughout and converges to a near-maximal reward. RecurrentPPO is intermittently competitive but prone to fluctuations. DQN and A2C both show moderate terminal performance, with A2C stabilizing late but still achieving a lower reward ceiling.
What’s more, Table II provides a numerical summary of success rate, evaluation return, average speed, $\Delta$TTCP, and decision-making time for each approach. In the unsignalized intersection scenario, our method attains the highest success rate (88%) while balancing speed and safety margins. For high-speed ramp merging, it achieves 91% success and outperforms the baselines in average return. Notably, A2C, despite having the highest speed, completely fails (0% success), demonstrating that overly aggressive driving sacrifices safety and thus overall performance. In four-lane adaptive cruise, our method reaches a perfect success rate (100%) alongside near-optimal speed and return. Although Dilu shows better speed and safety margins than traditional DRL methods, its extended reasoning time limits online deployment. Overall, our approach surpasses both conventional DRL algorithms and the LLM-based Dilu, underlining its effectiveness and robustness across diverse scenarios.
V-C2 Ablation Study
V-C3 Teacher-Student Comparison
V-D Case Analysis
TABLE II: Comparison of security, efficiency, and real-time test results of different methods in multiple scenarios.
@@ V-D Case Analysis
To further explore the decision-making process and behavior characteristics of the proposed model in actual scenarios, we selected three representative cases, as shown in Fig. 8.
In the unsignalized intersection shown in Fig. 8(a), after the agent approaches the stop line of the intersection in step 4, it actively slows down to give way to other vehicles that arrive at the intersection first until step 6; when trying to accelerate again in step 12, it promptly observes that the vehicle in the adjacent lane is about to pass, quickly judges and slows down again, and accelerates to leave after it passes. This behavior fully reflects the model’s understanding and compliance with traffic rules and social interactions, and can achieve safe and reasonable interactions with other traffic entities.
In the ramp merging scenario illustrated as Fig. 8(b), the model first accelerates quickly to complete the merging action; in step 6, it actively slows down to maintain a safe distance from the vehicle in front, and accelerates again after confirming that there is enough safe distance between it and the vehicle in front in step 14, and finally merges smoothly into the main road. This process shows that the agent has precise control over the acceleration and deceleration decisions in high-speed scenarios and a keen perception of risk factors.
In the example of four-lane adaptive cruise control shown in Fig. 8(c), the agent can continuously monitor and maintain a safe distance from the vehicle in front in dense traffic conditions or even in the presence of traffic disturbances, and adjust the speed in a timely manner to avoid rear-end collisions or excessive deceleration. This case shows that the model has good stability and active safety in long-term cruise tasks.
From the above cases, it can be seen that our model can demonstrate good interaction capabilities and strategic decision-making levels in a variety of complex driving scenarios, which further supports the conclusions of the aforementioned quantitative experiments.
VI Vehicle-in-Loop Experiment
VI-A Virtual-Real Integration Experimental Platform
To further assess the robustness and real-time performance of TeLL-Drive, we conduct a vehicle-in-loop experiment that combines virtual and real-world testing. A fusion platform is developed to integrate virtual traffic simulations with real vehicle hardware, allowing for the evaluation of the intelligent driving function in dynamic and complex traffic environments. This experimental setup enables the testing of autonomous driving decision-making under various conditions, including scenarios with potential safety hazards, both in virtual and real-world settings.
The virtual-real fusion platform consists of two main components: the AV hardware and traffic flow simulation software. As shown in Fig. 9, the traffic flow simulation software generates a virtual traffic environment, providing background traffic data that interacts with the real-world data captured by the AV’s sensors. These sensors collect real-time environmental information, which is then fused with the simulated traffic data through a data fusion process. This combined perception is transmitted to the planning control unit, which uses it to generate the vehicle’s motion trajectory. The resulting vehicle trajectory is then fed back into the simulation software, allowing for interaction between the AV and the virtual traffic flow.
Figure 9: The virtual-reality fusion experimental platform built based on Tongji Smart Town, which achieves complete synchronization between the virtual and real worlds. The autonomous driving vehicle uses the TeLL-Dirve decision algorithm, and the virtual vehicle uses TESSNG simulation operation.
In this experiment, the intelligent agent trained under the TeLL-Drive framework serves as the decision-making algorithm for the autonomous vehicle. The simulation vehicle operates using TESSNG[37], a high-level microscopic simulation software, which enables detailed modeling of vehicle dynamics in complex traffic scenarios. The experiment was conducted at Tongji University’s Autonomous Driving Smart Town, with the unprotected left turn at a complex intersection chosen as the test scenario. As the scenario with the lowest success rate in the simulation experiment, this scenario has inherent safety risks and requires precise decision-making in a dynamic environment. The integration of high-precision maps, precise timing positioning and full-element digitization enable complete synchronization between the real-world and virtual-world, ensuring that both environments are in sync during testing. This vehicle-in-loop setup provides a comprehensive platform for evaluating the performance of TeLL-Drive in real-time, dynamic driving scenarios.
VI-B Case Study Analysis
Figure 10: A real vehicle-in-loop experiment based on the virtual-reality fusion experimental platform. The vehicle equipped with TeLL-Drive choose to yield at the intersection in case 1 and has priority in case 2. In both scenarios, vehicles have complex social interactions.
We conducted real-vehicle experiments on the virtual-real fusion platform to evaluate the performance of the TeLL-Drive framework. Two representative cases are shown in Fig. 10, each captured from various perspectives, including the virtual twin platform perspective, the drone bird’s-eye view, the car-following view, the roadside view, and the in-car perspective. These multiple angles allow for a comprehensive analysis of the algorithm’s performance across different scenarios. The specific experimental video can be accessed on our website [^1] Vehicle-in-Loop Experimental Validation Video Weblink.
In Case 1, the autonomous vehicle equipped with TeLL-Drive begins from a standstill and accelerates toward the intersection. As it approaches the stop line, the vehicle slows down to create sufficient observation and decision space, enhancing its ability to assess the surrounding traffic. By the 7th second, the vehicle encounters an oncoming vehicle. Upon assessing the situation, the vehicle decides to slow further at the 12th second to yield and avoid a collision. After the oncoming vehicle passes, the autonomous vehicle resumes acceleration and approaches the exit road of the intersection by the 15th second. To maintain a safe distance from the vehicle in front, the system performs adaptive acceleration and deceleration, ensuring both safety and traffic efficiency. In this case, the autonomous vehicle is the last to leave the intersection, but the maneuver was executed safely and efficiently.
In Case 2, the autonomous vehicle follows similar actions up to the point before entering the intersection. However, at the 6th second, the vehicle observes fewer vehicles in the intersection and determines it can pass first, so it accelerates. By the 8th second, a vehicle on the left side approaches, prompting an interaction. After a brief period of strong interaction between the two vehicles, the simulation vehicle (SV) decides to slow down and stop, while our autonomous vehicle continues to pass first, successfully navigating the intersection.
These two cases demonstrate the robustness and reliability of the TeLL-Drive framework when deployed on real vehicles. The system effectively adapts to dynamic traffic scenarios, ensuring safe and efficient decision-making in complex environments. The ability of TeLL-Drive to handle both cooperative and conflict-driven interactions, while maintaining safety and traffic flow, underscores its potential for real-world autonomous driving applications.
VII Conclusion
@@
Our proposed TeLL-Drive framework integrates teacher-guided learning with attention-based policy optimization, enabling efficient knowledge transfer and robust decision-making. Experimental results demonstrate that TeLL-Drive outperforms conventional DRL methods and existing LLM-based approaches across multiple metrics, including success rate, average return, and real-time feasibility. Additionally, ablation studies highlight the significance of each model component, particularly the synergy between attention mechanisms and LLM teacher guidance. Finally, vehicle-in-the-loop experiments verify the robustness effectiveness of the model when deployed in practice. These findings confirm that our approach not only accelerates policy convergence but also enhances safety and adaptability across diverse traffic conditions. In the future, we will explore the application of the TeLL-Drive framework to more dynamic, multi-agent environments and verify its scalability and real-time adaptability through real-vehicle experiments in open road scenarios.